OpenStudy (anonymous):

|cosx-x| divide 3-|-xsquare| is an even or odd function?

OpenStudy (anonymous):

\(f(x) = \frac{|\cos x - x|}{3-|x^2|}\) \(f(-x) = \frac{|\cos (-x) + x|}{3-|(-x)^2|}=\frac{|\cos x + x|}{3-|x^2|}\) even means \(f(x)=f(-x)\), odd means \(f(x) \neq f(-x)\)